OpenStudy (anonymous):

What are the values of x for which 8-x^2 < 5? Explain.(:

OpenStudy (anonymous):

-x^2 < -3 x^2 > 3 x > sqrt{3} U x < -sqrt{3}

OpenStudy (anonymous):

Thanks.<3

OpenStudy (anonymous):

\[3-x^2<0\] is a start. this is a parabola that faces down, so it will be negative outside the zeros and positive between them. the zeros are \[\pm\sqrt{3}\] so your solution is waht pk51 wrote \[(-\infty, -\sqrt{3})\cup (\sqrt{3},\infty)\]